# Military AI Needs Technically-Informed Regulation to Safeguard AI Research and its Applications

### v1

@@ 3 Oversight Criteria for AI-LAWS, Grounded in System Behavior
The United States and the United Kingdom have introduced policies to govern military autonomy, but these efforts remain incomplete. While the U.S. Department of Defense’s Directive 3000.09 and the U.K.’s 2024 guidance outline principles for AI oversight, they do not specify how those principles apply to modern, adaptive AI-LAWS. Directive 3000.09 formally applies to all lethal autonomous weapon systems, including those that incorporate AI. It articulates high-level goals—such as ensuring AHJ, resilience to adversarial attacks, and review after model updates [59]. However, it offers little guidance on what these goals mean in practice—for example, it does not define what qualifies as a “substantial” model update requiring revalidation, or how often retrained or continually learning models should be re-evaluated. As a result, critics—including voices within the defense industry—warn that some AI-LAWS may enter deployment without rigorous review [44].
@@
The U.K. framework sets a stronger foundation. Its 2024 guidance proposes certification processes, third-party audits, and red-line prohibitions on delegating specific decisions to autonomous agents [55]. However, the proposal is not a binding regulatory framework, and key questions—such as how to validate adaptive models or define escalation-related boundaries—remain unresolved. The next section helps address this oversight gap, by offering specific initial policy proposals to codify.
@@ 4 Policy Recommendations for AI-LAWS Oversight
The following recommendations address urgent, technically grounded failure modes already visible in real-world military AI systems. While implementation rests with national and institutional actors, each recommendation highlights ways AI researchers can help define performance thresholds, anticipate failure modes, and ensure oversight reflects how these systems actually function.
@@
Of all AI-LAWS use cases, nuclear command and control is the clearest red line. Some argue that delegating or augmenting nuclear command with AI could improve deterrence over ambiguous or manually executed threats [77]. This view is extremely dangerous. Systems that delegate launch or targeting decisions in nuclear scenarios pose disproportionate risks—due to the stakes, the execution speed, and the unpredictability, opacity, and drift of AI under real-world conditions. Recent advances in AI speed and complexity have amplified the risk of escalation—via misclassification, drift, or adversarial manipulation [2, 14]. Yet tests of AI-based nuclear advisors date back decades [33].
We support formal, binding prohibitions on the use of AI systems in nuclear deployment pipelines. Existing proposals—including bilateral U.S.–China talks and a 2023 U.S. Senate bill [78, 79]—offer strong starting points, but lack international commitment and technical specificity. The 2022 National Defense Strategy committed the U.S. to require a human in the loop for any nuclear launch [80], but made no mention of AI advising systems. In 2023, the P3 (U.S., U.K., and France) issued more AI-specific declarations in the P5 council, urging China and Russia to follow suit [81].
Key message: Future policy should explicitly ban AI-based military systems from making nuclear launch decisions, regardless of human oversight. AI advising on nuclear strategy should either be banned outright or tightly constrained with technical specificity [2, 14, 43].
@@
Many risks from AI-enabled military systems stem from insufficient validation—such as poor generalization performance, adversarial vulnerability, and limited human interpretability. These challenges are well known in civilian AI [42], but their consequences are amplified in combat settings. Without shared standards for evaluating AI-LAWS under real-world and mission-specific constraints, oversight remains inconsistent and easily circumvented.
@@
We recommend developing international validation protocols tailored to AI-LAWS. Unlike past proposals that would have mandated fixed oversight models (e.g., hard human-in-the-loop controls for low-level decisions [70, 82]), we do not advocate a one-size-fits-all standard. Instead, we call for a coordinated international process to define what “sufficient oversight” means—anchored in technical performance rather than philosophical abstraction. This effort should include behavioral benchmarks, context-shift performance thresholds, and documentation of failure modes that may emerge only after deployment. Real-world performance should take precedence over interpretability, which remains poorly standardized for complex models [83, 84, 85, 86]. The goal is not to constrain architectures, but to ensure AI-LAWS reliably identify targets, adapt safely, and behave predictably under stress. Notably, setting such standards can be done without disclosing sensitive details on AI-LAWS, and individual governments would handle the actual validation process using their own personnel. Several countries and organizations have proposed promising components of such a framework [55, 59, 87, 53], which should now be formalized, stress-tested, and updated in collaboration with AI researchers. AI researchers can help shape criteria, vet claims, and clarify which risks remain invisible by default.
Because AI-LAWS technologies, deployments, and risks evolve rapidly, validation standards cannot be static. Any effective governance framework must be iterative—able to adapt as new behaviors emerge, failure modes are discovered, or architectures evolve. We recommend that participating governments coordinate through a standing international forum—not a fixed treaty—to maintain, revise, and share oversight standards. This approach supports both transparency and adaptability as the technical frontier advances.
@@
We propose a voluntary international consortium where participating states collaboratively define and refine AI-LAWS oversight standards. Membership in this consortium would grant representation in shaping criteria—ensuring buy-in, legitimacy, and flexibility across political and technical domains. Standards would evolve iteratively—based on new deployments, emerging failure modes, and evolving technologies—much like internet protocols or global financial risk auditing systems [88, 89, 90]. This model balances sovereignty with shared responsibility. It avoids the rigidity of treaty law while creating a durable mechanism for international norm-setting. In practice, the consortium could coordinate audits, verify compliance, and serve as a shared reference point for certifying AI safety in defense systems.
@@
Key message: Effective AI-LAWS oversight must be globally coordinated yet technically flexible. A standing international standards body—not a static treaty—is essential to ensure governance keeps pace with evolving capabilities.
@@
Recent proposals have explored using large language models and other generative AI systems to direct human combatants—a concept sometimes referred to as “minotaur warfare” [91, 92]. In this model, AI systems act as battlefield commanders—issuing orders, making strategic recommendations, or coordinating multiple units. While such approaches are technologically novel, they introduce significant risks with little demonstrated benefit.
Generative and decision-support AI systems often exhibit hallucinations, drift, adversarial vulnerability, and misplaced confidence [16, 93, 94, 95, 96, 97, 43, 42]. These characteristics are particularly dangerous in high-stakes, real-time decision contexts like military command [76, 98, 16, 14]. Equally concerning is growing evidence that humans tend to overtrust such systems—especially when they appear authoritative or fluent [99, 29]. In command settings, this risk can distort accountability, suppress dissent, and create institutional overreliance on opaque AI systems. This is not a call for Luddism. It is a recognition that delegating command decisions amplifies risks across all three domains: field-level failure, escalatory instability, and institutional erosion. We therefore recommend a categorical ban on AI systems that issue battlefield commands to human soldiers—whether through natural language interfaces or direct control platforms.
@@
As military systems increasingly rely on publicly available AI infrastructure—cloud compute, foundation models, and research personnel—the line between civilian and military assets is becoming dangerously blurred. Under Protocol I of the Geneva Conventions, civilian objects become lawful military targets if they make an effective contribution to military action. Yet it remains unclear how this doctrine applies to AI systems—such as models, training data, and compute infrastructure—used in both civilian and military contexts. Without clear guidance, a wide range of actors—including universities, commercial labs, and even individual researchers—could become valid military targets if their work is repurposed into AI-LAWS [100]. This creates unacceptable ambiguity for technical personnel who may unknowingly contribute to military systems. It also raises serious safety and ethical concerns for companies and institutions that host dual-use models or infrastructure. We recommend international treaties or norms that explicitly define when and how AI infrastructure—models, data, compute platforms, and personnel—qualifies as a military objective. This includes dual-use thresholds and legal boundaries for open-source contributions.
@@
The increasing convergence between civilian AI research and defense institutions—via funding, infrastructure, and personnel—is eroding long-standing boundaries between the two spheres [16, 15, 101, 39, 38]. This shift often occurs through formal programs such as dual-use labs or military contracting. Examples include proposals from the U.S. Defense Innovation Unit (DIU) to embed military AI research centers in universities [15], and China’s civil-military fusion strategy, which integrates academic and defense institutions [32, 102]. In both cases, research departments may become partially governed by national security interests, affecting what can be published, who can be hired, and how international partnerships function. While some collaboration is inevitable, blurred boundaries in historically civilian institutions risk undermining academic freedom, reducing international cooperation, and deterring researchers from working on civilian or humanitarian goals [30, 31, 12]. These shifts create a research ecosystem where technical expertise in AI is increasingly channeled toward military applications—sometimes without clear ethical guardrails or institutional independence. It may also introduce new AI safety risks, as rapid militarization of foundation models incentivizes premature deployment—especially in closed or classified settings [3, 43].
We recommend that institutions and companies publicly declare their own guidelines for preserving the independence of civilian AI research, since no one-size-fits-all policy can apply across sectors or nations. These declarations were once common, but many institutions have quietly moved away from them [103, 104, 105], creating confusion and concern among employees [106, 107]. Declarations should include disclosure of military funding and affiliations, and commit to separating civilian and military personnel, models, infrastructure, and funding—aligned with institutional goals and protections. Academic institutions should safeguard research agendas from redirection during reorganizations tied to defense contracting. They should also protect students and early-career researchers from being drawn into classified work without clear opt-in pathways. Review boards could be considered to assess the civil-military implications of AI projects—similar to existing models for biosafety and human subjects research. These structures can help ensure that dual-use concerns are addressed without restricting basic scientific inquiry.
@@ 5 Conclusion
Integrating ML algorithms into weapon systems has created AI-powered lethal autonomous weapons (AI-LAWS)—a class of system that behaves differently from conventional LAWS and demands new governance. These systems pose structural risks—such as unpredictability, opacity, and post-deployment adaptation—that existing oversight frameworks are not designed to address. Regulatory discussions have been slowed by vague definitions, abstract debates, and—most critically—a lack of engagement from AI experts. We introduced two criteria as a tractable basis for determining when AI-LAWS warrant oversight. They provide a foundation for targeted policy and validation standards. Our recommendations offer a practical starting point for AI-LAWS regulation, which is grounded in current and near future capabilities and aligned with technical foundations and realities. Realizing these proposals will require deeper engagement from AI researchers—especially in defining performance thresholds, stress-testing deployed systems, and shaping validation practices grounded in real-world behavior and the algorithms involved. Military AI governance is a deeply technical challenge that demands scientific input, and AI researchers are uniquely positioned to help ensure these systems are evaluated, constrained, and held accountable.

### v2

@@ 3 Oversight Criteria for AI-LAWS, Grounded in System Behavior
The United States and the United Kingdom have introduced policies to govern military autonomy, but these efforts remain incomplete. While the U.S. Department of Defense’s Directive 3000.09 and the U.K.’s 2024 guidance outline principles for AI oversight, they do not specify how those principles apply to modern, adaptive AI-LAWS. Directive 3000.09 formally applies to all LAWS, including those that incorporate AI. It articulates high-level goals—such as ensuring AHJ, resilience to adversarial attacks, and review after model updates [DoD2024]. However, it offers little guidance on what these goals mean in practice—for example, it does not define what qualifies as a “substantial” model update requiring revalidation, or how often retrained or continually learning models should be re-evaluated. As a result, critics—including voices within the defense industry—warn that some AI-LAWS may enter deployment without rigorous review [palantir2023].
@@
The U.K. framework sets a stronger foundation. Its 2024 guidance proposes certification processes, third-party audits, and red-line prohibitions on delegating specific decisions to autonomous agents [ukresponse2024]. However, the proposal is not a binding regulatory framework, and key questions—such as how to validate adaptive models or define escalation-related boundaries—remain unresolved. The next section helps address this oversight gap, by offering specific initial policy proposals to codify.
@@ 4 Policy Recommendations for AI-LAWS Oversight
The following recommendations address urgent, technically grounded failure modes already visible in real-world military AI systems. While implementation rests with national and institutional actors, each recommendation highlights an issue where AI researchers can help define performance thresholds, anticipate failure modes, and ensure oversight reflects how these systems actually function.
@@
Of all AI-LAWS use cases, nuclear command and control is the clearest red line. Some argue that delegating or augmenting nuclear command with AI could improve deterrence over ambiguous or manually executed threats [schwartz2025out, Hirsh2025-ak]. This view is extremely dangerous. Systems that delegate launch or targeting decisions in nuclear scenarios pose disproportionate risks—due to the stakes, the execution speed, and the unpredictability, opacity, and drift of AI under real-world conditions. Recent advances in AI speed and complexity have amplified the risk of escalation—via misclassification, drift, or adversarial manipulation [rivera2024escalation, atalan2025critical, souly2025poisoning]. Yet tests of AI-based nuclear advisors date back decades [Geist2016-kn].
We support formal, binding prohibitions on AI system control over nuclear deployment pipelines. Existing proposals—including bilateral U.S.–China talks and a 2023 U.S. Senate bill [renshaw2024biden, Markey2023-zy]—offer strong starting points, but lack international commitment and technical specificity. The 2022 National Defense Strategy committed the U.S. to require a human in the loop for any nuclear launch [posture2022], but made no mention of AI advising systems. In 2023, the P3 (U.S., U.K., and France) issued more AI-specific declarations in the P5 council, urging China and Russia to follow suit [Torode2024-sk].
Key message: Future policy should explicitly ban AI-based military systems from making nuclear launch decisions, regardless of human oversight. AI advising on nuclear strategy should either be banned outright or tightly constrained with technical specificity [rivera2024escalation, atalan2025critical, Biddle2024-ai].
@@
Many risks from AI military systems stem from insufficient validation—such as poor generalization performance, adversarial vulnerability, and low human interpretability. These challenges are well known in civilian AI [fodorbenchmark], but their consequences are amplified in combat settings. Without shared standards for evaluating AI-LAWS under real-world constraints, oversight remains inconsistent and easily circumvented.
@@
We recommend developing international validation protocols tailored to AI-LAWS. Unlike past proposals that would have mandated fixed oversight models (e.g., hard human-in-the-loop controls for low-level decisions [Article2014-vf, roff2016meaningful]), we do not advocate a one-size-fits-all standard. Instead, we call for a coordinated international process to define what “sufficient oversight” means—anchored in technical performance rather than philosophical abstraction. This effort should include behavioral benchmarks, context-shift performance thresholds, and documentation of failure modes that may emerge only after deployment. Real-world performance should take precedence over interpretability, which remains poorly standardized for complex models [siu2023stl, geirhos2023don, hurley2024stl, Hendrycks2025-lp]. The goal is not to constrain architectures, but to ensure AI-LAWS reliably identify targets, adapt safely, and behave predictably under stress. Notably, setting such standards can be done without disclosing sensitive details on AI-LAWS, and individual governments would handle the actual validation process using their own personnel. Several countries and organizations have proposed promising components of such a framework [ukresponse2024, DoD2024, ArmsControl2023, HD2021], which should now be formalized, stress-tested, and updated in collaboration with AI researchers. AI researchers can help shape criteria, vet claims, and clarify which risks remain invisible by default.
While it is not pragmatic nor desirable to encourage hostile nations to share knowledge for making more lethal and dangerous weapon systems, it is worth establishing an international forum or agency to develop evolving best practices in validating military AI. Indeed, there are already considerations in the UN for establishing a “global AGI observatory, certification system for secure and trustworthy AGI, a UN Convention on AGI, and an international AGI agency” for parallel concerns in the domain of hypothetical AI with superhuman intelligence [UN2025]. However, unlike AGI/ASI, dangerous and impactful military AI already exists, yet no specialized observatory, convention or agency is being considered to handle the AI-specific concerns around frontier military AI systems, if they do not meet the difficult-to-quantify threshold for AGI/ASI [Zeff2024-oy, hendrycks2025agidefinition].
@@
We propose a voluntary international consortium where participating states collaboratively define and refine AI-LAWS oversight standards. Membership in this consortium would grant representation in shaping criteria—ensuring buy-in, legitimacy, and flexibility across political and technical domains. Standards would evolve iteratively—based on new deployments, emerging failure modes, and evolving technologies—much like internet protocols, cybersecurity, or global financial risk auditing systems [mueller2007internet, van2013governance, romano2014diversity, efrony2018rule]. This model balances sovereignty concerns with the shared responsibility to rigorously manage the development of AI-LAWS given the potential risks. It also avoids the rigidity of treaty law while creating a durable mechanism for international norm-setting. In practice, the consortium could coordinate audits, verify compliance, and serve as a shared reference point for certifying AI safety in defense systems.
@@
Key message: Effective AI-LAWS oversight must be globally coordinated yet technically flexible. An international standards body is essential to ensure governance keeps pace with evolving capabilities.
@@
Recent proposals have explored using large language models and other generative AI systems to direct human combatants—a concept sometimes referred to as “minotaur warfare” [sparrow2023minotaurs, lushenko2024artificial]. In this model, AI systems act as battlefield commanders—issuing orders, making strategic recommendations, or coordinating multiple units. While such approaches are technologically novel, they introduce significant risks with little demonstrated benefit as the scale of the AI command increases and human oversight decreases. Some scholars argue that AI may actually impose a requirement for greater human oversight [goldfarb2021prediction].
Generative and decision-support AI systems often exhibit hallucinations, drift, adversarial vulnerability, and misplaced confidence [Lushe2024, yang2024poisoning, martinez2024re, dayan2024age, Jaz2025, Montii2025, Biddle2024-ai, fodorbenchmark]. These characteristics are particularly dangerous in high-stakes, real-time decision contexts like military command [Hunter2022-bw, johnson2022ai, Lushe2024, atalan2025critical]. Equally concerning is growing evidence that humans tend to overtrust such systems—especially when they appear authoritative, sycophantic or fluent [holbrook2024overtrust, buccinca2021trust, el2025moloch, sharma2023towards]. In command settings, this risk can distort accountability, suppress dissent, and create institutional overreliance on opaque AI systems. This is not a call for Luddism. It is a recognition that delegating command decisions amplifies risks across all three domains: field-level failure, escalatory instability, and institutional erosion. We therefore recommend a categorical ban on AI systems that autonomously command human soldiers with minimal human oversight—replacing human officers with AI officers.
@@
As military systems increasingly rely on publicly available AI infrastructure—cloud compute, foundation models, and research personnel—the line between civilian and military assets is becoming dangerously blurred. Under Protocol I of the Geneva Conventions, civilian objects become lawful military targets if they make an effective contribution to military action. Yet it remains unclear how this doctrine applies to AI systems—such as models, training data, and compute infrastructure—used in both civilian and military contexts. Without clear guidance, a wide range of actors—including universities, commercial labs, and even individual researchers—could become valid military targets if their work is repurposed into AI-LAWS [HoroLaw2024]. This creates unacceptable ambiguity for technical personnel who may unknowingly contribute to military systems. It also raises serious safety and ethical concerns for companies and institutions that host dual-use models or infrastructure. For example, one of the bombing targets in the brief 2025 Iran-Israel conflict was the Weizmann Institute of Science which houses frontier AI research labs [Kintisch2025-ok]. While it’s not known whether the AI labs were the primary motivation for the bombing, it’s suggestive that civilian research centers may be targeted more frequently as AI becomes more central to warfare. We recommend international treaties or norms that explicitly define when and how AI infrastructure—models, data, compute platforms, and personnel—qualifies as a military objective. This includes dual-use thresholds and legal boundaries for open-source contributions.
@@
The increasing convergence between civilian AI research and defense institutions—via funding, infrastructure, and personnel—is eroding long-standing boundaries between the two spheres [Lushe2024, shah2024unit, maaser2022big, ryseff2022exploring, Kreps2025-vm]. This shift often occurs through formal programs such as dual-use labs or military contracting. Examples include proposals from the U.S. Defense Innovation Unit (DIU) to embed military AI research centers in universities [shah2024unit], and China’s civil-military fusion strategy, which integrates academic and defense institutions [ASPI_2025, Margarita_Konaev2023-qv]. In both cases, research departments may become partially governed by national security interests, affecting what can be published, who can be hired, and how international partnerships function. While some collaboration is inevitable, blurred boundaries in historically civilian institutions risk undermining academic freedom, reducing international cooperation, and deterring researchers from working on civilian or humanitarian goals [Daniels2019-ib, kania2021myths, simmons2024ai]. These shifts create a research ecosystem where technical expertise in AI is increasingly channeled toward military applications—sometimes without clear ethical guardrails or institutional independence. It may also introduce new AI safety risks, as rapid militarization of foundation models incentivizes premature deployment—especially in closed or classified settings [caballero2024large, Biddle2024-ai].
We recommend that institutions and companies publicly declare their own guidelines for preserving the independence of civilian AI research, since no one-size-fits-all policy can apply across sectors or nations. These declarations were once common, but many institutions have quietly moved away from them [Biddle2024-fz, BD2023-nq, Tiku2025-tj], creating confusion and concern among employees [Perrigo2024-le, De-Vynck2024-xi]. Declarations should include disclosure of military funding and affiliations, and clarifications of any commitments to separating civilian and military personnel, models, infrastructure, and funding—aligned with institutional goals and protections. Academic institutions should safeguard research agendas from redirection during reorganizations tied to defense contracting. They should also protect students and early-career researchers from being drawn into classified work without clear opt-in pathways. Review boards could be considered to assess the civil-military implications of AI projects—similar to existing models for biosafety and human subjects research. These structures can help ensure that dual-use concerns are addressed without restricting scientific inquiry or leading to unacceptable risks.
@@
5 AI researchers can contribute on military AI policy
It is important to note that there are a number of means for academic and civilian researchers to contribute to military AI regulation, and that a career in policy or national security work is not required to be involved. Ways for AI researchers to get involved in these discussions include, but are not limited to: (1) Participate in international conferences or dialogues on AI safety and arms control [jon2024scientific, Rosen2024-md, hraim2024, idais]. (2) Publish research on AI red teaming and failure modes in contexts relevant to the military [rivera2024escalation]. (3) Collaborate with researchers at think tanks, policy institutes or military colleges which do public-facing policy research relevant to military AI [WP2025-gp, Reynolds2025-od]. (4) Encourage universities and departments to establish written norms and policies defining boundaries for military AI research, as institutions in the U.S. and Europe generally lack explicit guidelines on acceptable defense-related topics [Roberts2005, DAgostino2024-pw, Alviz2025-rz]. (5) Avoid overstating benchmark results. Researchers should recognize that both think tanks and military analysts closely monitor technical AI publications. It is important that our community does not overstate benchmark success as implying immediate readiness for complex real world applications, which could encourage rushed and risky adoption by non-experts [fodorbenchmark, gu2025illusion, Stone2025-rh].
@@ 6 Conclusion
Integrating frontier AI algorithms into weapon systems has created AI-LAWS—a class of system that behaves differently from conventional LAWS and demands new governance. These systems pose structural risks—such as unpredictability, opacity, and post-deployment adaptation—that existing oversight frameworks are not designed to address. Regulatory discussions have been slowed by vague definitions, abstract debates, and—most critically—a lack of engagement from AI experts. We introduced two criteria as a tractable basis for determining when AI-LAWS warrant oversight. They provide a foundation for targeted policy and validation standards. Our recommendations offer a practical starting point for AI-LAWS regulation, which is grounded in current and near future capabilities and aligned with technical foundations and realities. Realizing these proposals will require deeper engagement from AI researchers—especially in defining performance thresholds, stress-testing deployed systems, shaping validation practices grounded in real-world behavior and establishing AI safety principles for the military domain. Military AI governance is a deeply technical challenge that demands scientific input, and AI researchers are uniquely positioned to help ensure these systems are evaluated, constrained, and held accountable.